OpenStudy (anonymous):

Gentry poured 23.5 ounces of milk and 7.75 ounces of cream in a mixing bowl. How much more cream could Gentry add to the bowl if the capacity of the bowl is 35.5 ounces?

OpenStudy (anonymous):

23.5 + 7.75 =?

OpenStudy (anonymous):

the answer is 4.25

OpenStudy (anonymous):

23 + 7 = 30. .75 +.50 = 1.25 30+1.25= 31.25

OpenStudy (tkhunny):

This may not be a simple question, due to molecular spacing, but let's assume it's as simple as it looks. Better yet, 35.5 - 23.5 - 7.75 = ?? Why are you struggling with this? Where were you? What had you done?

OpenStudy (anonymous):

35.5-31.25=4.25

OpenStudy (anonymous):

thanks that's what I thought I needed to make sure

OpenStudy (tkhunny):

Here's the plan... Next time, show us what you thought up front, in the very first post on a new thread.